Team updates before Christmas


The weather this week has kept things fairly stable, with temperatures hovering right around the freezing mark. There were katabatic winds in the valleys on Monday that flattened several tents and one tent at Bonney blew away. There were snow squalls on Friday in the valleys temporarily delaying helicopter operations and slowing stream flow.

Mike Jackson left the ice on Saturday and Julie Palais takes over as the NSF Science rep until Jan 7.  The Christmas holiday is approaching and the galley has been decorated. Work centers including helo ops will be closed on Friday and Saturday, but Sunday will be a normal work day. 

•  C-504 (glaciers) Anna returned to McMurdo this week to process the second round of ice cores.  She has 21 cores to cut, melt and filter for particulates. Forrest stayed in the field to assist the artist, Helen Glazer, and to continue collecting melt water on the Canada Glacier. 

•  C-505 (limno) moved to Lake Bonney on Tuesday on schedule. They completed the sampling of the west lobe on Saturday. They are working on melting the hole to deploy the second sediment trap early next week.  The NSF Science rep approved sending the Lake Hoare assistant camp manager to oversee deployment of the sediment trap. 

•  C-506 (streams) The stream team has been busy sampling during the relatively high flows for mid-summer.  In Wright Valley they replace a nitrogen cylinder at the Vanda gauge and also visited the gauges in the Miers Valley at Miers and Adams Streams.  They traveled to Lake Bonney on Friday and to Lake Hoare on foot on Saturday. 

•  C-509 (geochem) has been analyzing water samples on the IC. I continue to work on setting up the new Shimadzu instrument in lab 107 with help from Nik. This has been a fairly slow but steady process including a push to clear out the remaining old stuff left behind by the contractor to make room for the new instrument and related supplies. 

• C-511 (met + lake physical) The team completed their final day trip this week and flew north schedule on Saturday.

• C-517 is also in the dry valleys.  Dave Van Horn is conducting planned soil experiments.  

The Good

•  Overall, things are going well. There is a bit of a lull in science in McMurdo as many teams finish their early season work and head north.  Package mail and freshies are trickling in with the reduced airlift, but things are moving.  Intercontinental flights have been on schedule. 

The Bad-

            We are about to start the 10-day gap in intercontinental flights.

The Ugly-

            Nothing to report.  

Article Section: